11. Suppose a car travels at a constant rate of 70 kph for 2 hours and travels 55 kph
thereafter. Show that distance travelled is a function of time and find the rule of the
function.
12. A man jogs for 20 minutes at a rate of 6 kph, then walks for 30 minutes at a rate of 3 kph,
then sits and rests for 15 minutes and finally walks for 45 minutes. Find the rule of the
function that expresses his distance travelled as a function of time
1
Expert's answer
2021-10-08T08:05:36-0400

11. "Distance=Speed \times Time"

Let "d(t)" be the distance function. Where "t" is the time.

"d(t)=70\times 2+55\times t\\
d(t)=55t+140"

It is clear that distance is a function of time "t" .


12. For the first 20 minutes, "d(t)=6t" when "0\leq t \le\frac{1}{3}"

Next 30 minutes, "d(t)=2+3(t-\frac{1}{3})" when "\frac{1}{3}<t\leq \frac{5}{6}"

Next 15 minutes, "d(t)=3.5" when "\frac{5}{6}< t\leq \frac{13}{12}"

Last 45 minutes, "d(t)=3.5+3(t-\frac{13}{12})" when "\frac{13}{12}<t\leq \frac{11}{6}"

Therefore,

"d(t)=\begin{cases}
6t ~~~~~~~~~~~~~~~~~~~~~~~~~~\text{ when } 0\leq t \leq\frac{1}{3}\\
2+3(t-\frac{1}{3}) ~~~~~~~~\text{ when } \frac{1}{3}<t\leq \frac{5}{6}\\\frac{5}{6}
3.5~~~~~~~~~~~~~~~~~~~~~~~~\text{ when } \frac{5}{6}< t\leq \frac{13}{12}\\
3.5+3(t-\frac{13}{12}) ~~~~~\text{ when } \frac{13}{12}<t\leq \frac{11}{6}


\end{cases}"